OpenStudy (anonymous):

Y=3-logx How is this done??

OpenStudy (unklerhaukus):

Is that \[y = \log_3(x)\]? use the log identity \[y = \log_b(x)\qquad\iff\qquad b^y=x \]

OpenStudy (anonymous):

Yeah I know that but how is it done with a negative between 3 and logx

OpenStudy (anonymous):

the answer is y=10^3-y but how

OpenStudy (unklerhaukus):

oh it is a negative? like this \[y = 3-\log(x)\]?

OpenStudy (unklerhaukus):

sorry for my earlier misinterpretation.

OpenStudy (anonymous):

No its suppose to be a negative between the three and y

OpenStudy (unklerhaukus):

oh , right i erred \[y = 3 - \log (x)\\ \log(x) = 3-y\\ x = 10^{3-y}\]

OpenStudy (unklerhaukus):

Is this what you were looking for?

OpenStudy (anonymous):

Wait so what exactly did you do?

OpenStudy (unklerhaukus):

\[y = 3 - \log (x)\]add log x to both sides of the equation, \[y+\log(x) = 3\]take away y from both sides .\[\log(x) = 3-y\]then applied log identity \[x = 10^{3-y}\]

OpenStudy (anonymous):

No I think I got it now! Thank you!